### v1

# Zero-Shot Machine-Generated Text Detection Using Mixture of Large Language Models

Abstract
The dissemination of Large Language Models (LLMs), trained at scale, and endowed with powerful text-generating abilities has vastly increased the threats posed by generative AI technologies by reducing the cost of producing harmful, toxic, faked or forged content. In response, various proposals have been made to automatically discriminate artificially generated from human-written texts, typically framing the problem as a classification problem. Most approaches evaluate an input document by a well-chosen detector LLM, assuming that low-perplexity scores reliably signal machine-made content. As using one single detector can induce brittleness of performance, we instead consider several and derive a new, theoretically grounded approach to combine their respective strengths. Our experiments, using a variety of generator LLMs, suggest that our method effectively increases the robustness of detection.
1 Introduction
Large Language Models (LLMs) have greatly improved the fluency and diversity of machine-generated texts. The release of ChatGPT and GPT-4 by OpenAI has sparked global discussions on the effective use of AI-based writing assistants. However, this progress has also introduced considerable threats such as fake news Zellers et al. (2019), and the potential for harmful outputs such as toxic or dishonest speech (Crothers et al., 2023), among others. As it seems, the research on methods aimed at detecting the origin of a given text to mitigate the dissemination of forged content and to prevent technology-aided plagiarism still lags behind the rapid advancement of AI itself. [^1] As illustrated by the discontinuation of OpenAI’s own detector https://openai.com/index/new-ai-classifier-for-indicating-ai-written-text/.
Many works have focused on tools that could spot such AI-generated outputs and address these underlying risks. From a bird’s eye view, this typically involves using detector models to discriminate generator models’ outputs from legitimate human writings. Multiple versions of this generic text classification task have been considered, varying e.g. the number of possible outcomes and the amount of available supervision (see Section 5). Owing to its large user base and applications, the largest effort has focused on one specific generator, ChatGPT, for which training and test data is easily obtained. Yet, the corresponding supervised binary problem, with a unique known generator, is not the only way to frame this task. A more challenging problem, that we study here, is generator-agnostic artificial text detection, where the models to detect are not known in advance.
As pointed out e.g. in Antoun et al. (2024); Hans et al. (2024); Wang et al. (2024a), the performance of artificial text detection systems varies depending on the choices of the detector(s) / generator(s) pair. This means, in principle, that the search for optimal detection performance should include a systematic exploration of the space of possible detectors. As the number and diversity of LLMs keep increasing, such exploration seems not only challenging but also unrealistic. Furthermore, Dugan et al. (2024) demonstrated shown that the current detection methods are easily fooled when changing the generator or the sampling method. Our proposal thus considers ensemble methods, where a coalition of several available models can be exploited to build the detector. This approach eliminates the need to empirically search for the best detector(s), and yields detection systems that can robustly detect multiple generators. Furthermore, it lends itself to the smooth extension of the ensemble as new models become available while improving generalization to unseen models.
Our main contribution. In this paper, using fundamental information-theoretic principles from universal compression, we derive a new ensemble method that optimally combines the strength of multiple LLMs . Relying on existing benchmarks, as well as a new corpus of machine-generated texts, we confirm in our experiments that this method yields detectors that can robustly detect a multiplicity of generators and that it compares favorably with several recent proposals. Our analyses explore the effect of adding models into the ensemble, and also highlight, using a multilingual dataset, how the contribution of each constituent model changes when the source of artificial texts is modified.
2 Robust Scoring Algorithms (RSA)
2.1 Background
We consider models for language generation tasks that define a probability distribution over strings. Formally, these language models are probability distributions over an output space $\mathcal{Y}$ which contains all possible strings over vocabulary $\Omega$: $\mathcal{Y}\triangleq\big{\{}\texttt{BOS}\circ\mathbf{y}\circ\texttt{EOS}\,|\,\mathbf{y}\in\Omega^{*}\big{\}},$ BOS and EOS denote for special reserved beginning-of-sequence and end-of-sequence tokens, respectively, and $\Omega^{*}$ denotes the Kleene closure of $\Omega$.
Today’s models for language generation are typically parameterized by encoder-decoder or decoder-only architectures with attention mechanisms with trainable weights $\theta\in\Theta$. These models follow a local-normalization scheme, meaning that $\forall$ $t>0$, $p_{\theta}(\cdot|\mathbf{y}_{<t},)$ defines a conditional probability distribution over $\bar{\mathcal{Y}}=\mathcal{Y}\cup\texttt{EOS}$. The probability of a sequence $\mathbf{y}=\langle y_{0},y_{1},\ldots\rangle$ is expressed as:
$p_{\theta}(\mathbf{y})=\prod_{t=1}^{T}p_{\theta}(y_{t}|\mathbf{y}_{<t}),$ (1)
and $\mathbf{y}_{<t}=\langle y_{0},\ldots,y_{t-1}\rangle$, with $y_{0}=\texttt{BOS}$.
Measurement of information. The fundamental concept in information is “surprisal”, using the relationship: $\textrm{information}=-\log(\textrm{probability})$  Cover and Thomas (2006), and assuming the use of coding techniques such as Huffman and Arithmetic codes Shields (1996) which give message lengths closely approximating the ideal length in binary digits. Here, the measure of information is most conveniently introduced in the context of lossless compression. That is, we will look a what happens when information is passed from a encoder to a decoder. Humans often use codes, such as natural languages, which are not optimal for any set of prior expectations. There are good reasons for natural languages to be less than “optimal” when regarded as codes for data compression. One reason is that spoken language is transmitted from speaker to listener via a noisy channel. Codes with substantial redundancy can tolerate some degree of corruption without becoming unintelligible, whereas optimal compression codes, in which every digit matters, are very sensitive to corruption. Artificial codes for information storage are often designed so that legal strings conform to a strict pattern while most strings do not. A corrupted received string can then be corrected by replacing it with the nearest legal string, provided the degree of corruption is not too great. However, for our purposes, we need not consider errors in the storage of messages, and hence need not be concerned with error-correcting codes.
Explanations of data. Given a body of text represented in a finite string $\mathbf{y}_{<t}=\langle y_{0},\ldots,y_{t-1}\rangle$, an “explanation” of the next token $y_{t}$ is a binary string encoding the symbol in a particular format with minimum length $\mathcal{L}_{\theta}(y_{t}|\mathbf{y}_{<t})\triangleq-\log p_{\theta}(y_{t}|\mathbf{y}_{<t})$. Its expected value is termed conditional entropy:
$\mathcal{H}_{\theta}(Y_{t}|\mathbf{y}_{<t})=\sum_{y_{t}\in\Omega}p_{\theta}(y_{t}|\mathbf{y}_{<t})\mathcal{L}_{\theta}(y_{t}|\mathbf{y}_{<t}).$
Finally, another important concept is the conditional mutual information (MI) between two random variables $\mathbb{M}$ and $Y_{t}$, given a sequence value $\mathbf{y}_{<t}$, defined as Cover and Thomas (2006):
$\mathcal{I}_{\theta}(\mathbb{M};Y_{t}|\mathbf{y}_{<t})=\mathcal{H}_{\theta}(Y_{t}|\mathbf{y}_{<t})-\mathcal{H}_{\theta}(Y_{t}|\mathbb{M},\mathbf{y}_{<t}),$
$\mathcal{H}_{\theta}(Y_{t}|\mathbb{M},\mathbf{y}_{<t})=\underset{m\sim\mu(m|\mathbf{y}_{<t})}{\mathbb{E}}\left[\mathcal{H}_{\theta}(Y_{t}|m,\mathbf{y}_{<t})\right].$
It captures the amount of information we get about $\mathbb{M}$ when observing $Y_{t}$, and already knowing $\mathbf{y}_{<t}$.
2.2 Combining LLMs with Robust Scoring Algorithms (RSA)
Let $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})\triangleq\big{\{}p_{\theta}(\mathbf{y}|m)\,:\,m\in\mathcal{M}\big{\}}$ be a family of LLMs, as presented in (1), with identifying set of indexes =$\mathcal{M}=\{1,\dots,M\}$. Given $\mathbf{m}$ a sequence of $T$ indexes in $\mathcal{M}$, where $m_{t}$ specifies the model index for generating token $y_{t}$, we derive:
$p_{\theta}(\mathbf{y}|\mathbf{m})=\prod_{t=1}^{T}p_{\theta}(y_{t}|m_{t},\mathbf{y}_{<t}).$ (2)
Depending on the choice of explanation $m_{t}$ for token $t$, certain tokens in $\Omega$ become unsurprising (high probability) while others become very surprising or unbelievable (low or zero probability).
A family of LLMs can be exploited to produce explanations of token sequences. To this end, we assume some $\hat{m}_{t}=f_{t}(\mathbf{y}_{<t})$ which selects a probability distribution $p_{\theta}(y_{t}|\hat{m}_{t},\mathbf{y}_{<t})$ over $\Omega$. Given $\hat{m}_{t}$, the encoder can construct an optimum code for token $y_{t}$, using distribution $p_{\theta}(y_{t}|\hat{m}_{t},\mathbf{y}_{<t})$. Thus, a rich family of LLMs therefore allows us to capture and represent regular patterns in token sequences via the model selector $\hat{m}_{t}$ and subsequently use it to minimize the total expected codelength.
Identifying explanations of data. We now turn to the problem of determining an adequate sequence of models $\hat{\mathbf{m}}=\langle\hat{m}_{0},\ldots,\hat{m}_{T}\rangle$. Our goal will be to derive a robust scoring algorithm that best extracts regularity in the data, which is equivalent to identifying the model that achieves the best compression of the input tokens. Suppose we are given a family of LLMs $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$, with corresponding Shannon codelengths $\mathcal{L}_{\theta}(y_{t}|m,\mathbf{y}_{<t})\triangleq-\log p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$ for each $y_{t}$. These can be viewed as a collection of data compressors, indexed by $m$. We can measure the performance of encoding $y_{t}$ at time $t$ relative to $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$. If we chose to encode the token $y_{t}$ using a model $q(y_{t}|\mathbf{y}_{<t})$, the resulting expected excess codelength (or overhead) w.r.t. any distribution $p_{\theta}\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ is given by:
$\mathcal{R}_{\theta}(m,q\,;\mathbf{y}_{<t})\triangleq\underset{y_{t}\sim p_{\theta}(y_{t}|m,\mathbf{y}_{<t})}{\mathbb{E}}\Big{[}-\log q(y_{t}|\mathbf{y}_{<t})\Big{]}-\mathcal{H}_{\theta}(Y_{t}|m,\mathbf{y}_{<t})\vspace{1mm}$
which is non-negative since $\mathcal{H}_{\theta}(Y_{t}|m,\mathbf{y}_{<t})$ is the minimum expected codelength. $\mathcal{R}_{\theta}$ represents the extra averaged number of bits needed to encode $y_{t}$ using the code/LLM $q(y_{t}|\mathbf{y}_{<t})$, as compared to $\mathcal{H}_{\theta}(Y_{t}|m,\mathbf{y}_{<t})$, the number of bits needed if we would have used the best fitting LLM in $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ with hindsight. However, the encoder cannot know the underlying model artificially generating $y_{t}$ so we take a worst-case approach and look for universal LLMs with small worst-case expected overhead, where the worst-case is over all models in $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$. $\mathcal{R}_{\theta}$ is our quality measure and hence, the ‘optimal’ LLM relative to $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$, for a given context $\mathbf{y}_{<t}$, is the distribution minimizing:
$q^{\star}(y_{t}|\mathbf{y}_{<t})\triangleq\operatorname*{argmin}_{q\in\mathcal{P}(\Omega)}\,\max_{m\in\mathcal{M}}\mathcal{R}_{\theta}(m,q\,;\mathbf{y}_{<t}),\vspace{1mm}$ (3)
where the minimum is over all distributions on $\Omega$. The minimizer is the code with the smallest overhead (extra number of bits) compared to the optimal code that is best in hindsight in the worst-averaged case over all LLMs in $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$.
Leveraging codelengths for identifying AI-generated text. The averaged overhead of the optimal codelength $-\log q^{\star}(y_{t}|\mathbf{y}_{<t})$ obtained by solving (3) seems to be a very reasonable choice for building a robust score function to detect AI-generated text because of the following properties:
  • The better the best-fitting LLM in $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ fits the artificially generated data, the shorter the codelengh $\mathcal{L}^{\star}(y_{t}|\mathbf{y}_{<t})\triangleq-\log q^{\star}(y_{t}|\mathbf{y}_{<t})$.
  • No LLM in $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ is given a prior preference over any other since $\mathcal{R}_{\theta}(m,q^{\star}\,;\mathbf{y}_{<t})$ $\leq\mathcal{R}_{\theta}(m,p_{\theta}\,;\mathbf{y}_{<t})$ for all $p_{\theta}\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$, i.e., we are treating all LLMs within our universe $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ on the same footing.
These observations lead to the following score.
Definition 1 (Robust Scoring Algorithm).
The Robust Scoring Algorithm (RSA) $S_{\textrm{Av}}(\mathbf{w})$ for an input sentence $\mathbf{w}=\langle w_{0},w_{1},\ldots\rangle$ is defined as:
$S_{\textrm{Av}}(\mathbf{w})\triangleq\frac{1}{T}\displaystyle\sum_{t=1}^{T}\Big{[}\underbrace{\sum_{y_{t}\in\Omega}\mathbbm{1}[y_{t}=w_{t}]\mathcal{L}^{\star}(y_{t}|\mathbf{w}_{<t})}_{\textrm{(codelength for observed token)}}-\frac{1}{M}\sum_{m\in\mathcal{M}}\underbrace{\Big{(}\sum_{y_{t}\in\Omega}p_{\theta}(y_{t}|m,\mathbf{w}_{<t})\mathcal{L}^{\star}(y_{t}|\mathbf{w}_{<t})\Big{)}}_{\textrm{(codelength for generated token using model `m')}}\Big{]},$
where $\mathcal{L}^{\star}(y_{t}|\mathbf{w}_{<t})\triangleq-\log q^{\star}(y_{t}|\mathbf{w}_{<t})$. For a suitable $\delta>0$, if $S_{\textrm{Av}}(\mathbf{w})\geq\delta$, then the text $\mathbf{w}$ is declared to be human and otherwise AI-generated.
Remark 1.
The first term in $S_{\textrm{Av}}(\mathbf{w})$ represents the averaged per-token codelength of the input sequence for the code/LLM $q^{\star}$, which corresponds to the well-known perplexity. The second term is the averaged per-token codelength over all randomly generated sequences according to the averaged LLMs in $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$, which is the average of the cross-entropy with respect to all models in the family. The resulting score is the difference between these codelengths. If the input sentence is generated by one of the LLMs in the family or another closely related one, the score is expected to be small, as $q^{\star}$’s goal is to extract as much regularity as possible from $\mathbf{w}$. However, if the input sentence is human-generated, the score is expected to be large as the first term will dominate.
Remark 2.
An alternative score, denoted by ${S}_{\max}(\mathbf{w})$, can be defined by replacing the sum over $m$ with its maximum. ${S}_{\textrm{Av}}(\mathbf{w})$ and ${S}_{\max}(\mathbf{w})$ will be empirically evaluated.
The next proposition provides a theoretically result together with an efficient iterative algorithm to optimally solve expression (3). The proof of this proposition is relegated to Appendix A.
Proposition 1 (Optimal codelength).
The optimal solution to (3) is shown to be the mixture of LLMs:
$q^{\star}(y_{t}|\mathbf{y}_{<t})=\sum\limits_{m\in\mathcal{M}}\mu^{\star}(m|\mathbf{y}_{<t})p_{\theta}(y_{t}|m,\mathbf{y}_{<t}),$
where the distribution $\mu^{\star}(\cdot|\mathbf{y}_{<t})$ of the random variable $\mathbb{M}$ of LLM index on $\mathcal{M}$ follows by solving:
$\mu^{\star}(\cdot|\mathbf{y}_{<t})\triangleq\operatorname*{argmax}_{\mu\in\mathcal{P}(\Omega)}\,\mathcal{I}_{\theta}\big{(}\mathbb{M};Y_{t}|\mathbf{y}_{<t}\big{)}.$ (4)
Furthermore, the weights $\{\mu^{\star}(m|\mathbf{y}_{<t})\}_{m\in\mathcal{M}}$ can be computed efficiently by using Blahut–Arimoto algorithm, referred to as Blahut–Arimoto weights.
2.3 Implementation
Proposition 1 implies that in order to implement the scoring function introduced in Definition 1, it would be enough to solve the optimization in (4) which is much simpler than (3). Interestingly, Blahut–Arimoto algorithm (Arimoto, 1972; Blahut, 1972) provides us with an efficient iterative method to to compute the maximization of mutual information (4) (see Appendix B). This algorithm lies at the core of our scoring procedure.
Algorithm 1 RSA Scoring
1:Input: text $\mathbf{w}=\langle w_{0},w_{1},\ldots\rangle$, LLMs $(1,2,\ldots,M)$
2:for $w_{t}$ in $\mathbf{w}$ do
3:     $\mu^{*}(m|\mathbf{w}_{<t})\leftarrow\text{Blahut\textendash Arimoto }(\mathcal{P}_{\mathcal{M}}(\mathcal{Y});\mathbf{w}_{<t})$
4:     $q^{*}(y_{t}|\mathbf{w}_{<t})\leftarrow\sum_{m\in\mathcal{M}}\mu^{*}(m|\mathbf{w}_{<t})p_{\theta}(y_{t}|m,\mathbf{w}_{<t})\!\!\!$
5:     $s_{t}(\mathbf{w})\leftarrow\mathcal{L}^{\star}(w_{t}|\mathbf{w}_{<t})-\frac{1}{M}\sum_{m\in\mathcal{M}}\Big{(}\underset{y_{t}\sim p_{\theta}(y_{t}|m,\mathbf{w}_{<t})}{\mathbb{E}}\left[\mathcal{L}^{\star}(y_{t}|\mathbf{w}_{<t})\right]\Big{)}$
6:end for
7:$S(\mathbf{w})\leftarrow\frac{1}{T}\sum_{t}s_{t}(\mathbf{w})$triangleright RSA score over the whole text
3 Experimental Settings
3.1 Datasets & metrics
To evaluate our method on a diverse set of texts and generative models, we collected four publicly available datasets used in previous publications: Ghostbuster (Verma et al., 2023), Binoculars (Hans et al., 2024), M4 (Wang et al., 2024a) and a corpus of scholarly texts Liyanage et al. (2022).
  • The Ghostbuster dataset is split into three parts: WritingPrompts, based on the r/WritingPrompts subreddit where users submit stories in response to short prompts; Reuters, using the Reuters 50-50 authorship identification dataset (Houvardas and Stamatatos, 2006); and Essays, comprised of essays scraped from IvyPanda, [^2] https://ivypanda.com/ a website dedicated to homework help. Each part contains 1,000 original texts that have been regenerated using headlines with ChatGPT (using 5 different prompts) and with Claude (Anthropic, 2023) (one prompt), for a total of 7,000 texts.
  • The Binoculars dataset contains samples of human-written texts from CCNews, Pubmed and CNN; alternative completions are automatically generated using Llama-2-13b (Touvron et al., 2023) and Falcon-7b (Almazrouei et al., 2023). Their generation technique collects the first 50 tokens for each text sample and uses them as a prompt to generate machine output. Those first 50 tokens are then removed from the result so that samples only contain machine-generated texts.
  • The M4333For Multi-Lingual, Multi-Domain, Multi-Generator Machine-Generated text detection. paper describes a massive dataset of natural texts collected from a diverse set of sources. Comparable artificial texts are obtained by generating with 6 LLMs, with prompts such as article titles, headlines, or abstracts depending on their domain. In our experiments, we only use the two “multilingual” generators (ChatGPT  [^4] https://chatgpt.com/ and Davinci003555https://platform.openai.com/docs/models), and the balanced sets of $3,000$ pairs of (artificial, natural) texts in Chinese, Russian, Bulgarian, Arabic, and German.
  • The academic benchmark proposed by Liyanage et al. (2022) is generated using a GPT2 model Radford et al. (2019) fine-tuned on papers from Arxiv’s “Computation and Language" (CS.CL) domain. 100 seed texts are used to generate new papers of similar length, using the first 50 words as prompts. Sections such as “methodology”, “results”, “evaluation”, and “discussion” are voluntarily omitted from texts, to ensure that discriminating factors do not include the comparative use of diagrams, tables, and equations.
Table 1 displays the main statistics for these 4 test sets. For completeness, we also augmented the Binoculars dataset using Llama-2-7b as an alternative text generator with Huggingface’s transformers model.generate() (Wolf et al., 2020). [^6] https://huggingface.co/docs/transformers. Sampling: repetition_penalty: 1.18, temp.: 0.6, top_p: 0.9. As for the original corpus, the first 50 tokens of the original texts provide the starting context for both machine & human texts. To test our method on extreme cases, we also generated 3000 texts of 500-token length using a unigram model trained on the Brown Corpus (Francis and Kucera, 1979), along with original extracts of the same length for comparison, details are in section  4.1.4. These new corpora will be released with our detector’s implementation.
Human Artificial
Corpus Name Generator(s) # texts avg length # texts avg length
Binoculars Llama-2-13b, Falcon 9,148 2,252 11,178 677
Binoculars+ Llama-2-7b, Mistral-7b 9,148 2,202 18,295 397
Ghostbuster ChatGPT (×$\times{}5$), Claude 1,000 826 6,000 754
arxiv-cs.cl GPT2 100 1,940 100 1977
M4 (Multilingual) ChatGPT, Davinci-003. 24,000 729 24,000 649
Table 1: Human [H] and Artificial [A] texts used in the experiments, all lengths are in Llama-2 tokens.
Taken together, these datasets represent a large variety of genres, styles, themes, languages, and generators, allowing us to thoroughly assess the strengths and weaknesses of our detection strategy.
Metrics.
As in most studies, we report the AUROC score as our main evaluation metric. Depending on the application, True Positive Rate (TPR) for a predefined False Positive rate (e.g., 5%) is also worth looking at and is also reported in most of our results. It is important to note that we compare imbalanced datasets in the case of Ghostbuster and the original Binoculars datasets, see numbers in Table  1, whereas, for our Binoculars regenerations and the M4 multilingual datasets, the data is comprised of (original, generated) pairs of texts.
3.2 Baselines
Machine-generated text detection methods are usually divided into two main categories (Dugan et al., 2024): supervised and unsupervised (metric-based). The former uses supervision data to fine-tune a pre-trained model for the detection task, typically focusing on some known generators. Our method belongs to the latter family, as our metric (the RSA score of 2) does not require any training data.
Accordingly, we compare our method to other zero-shot unsupervised techniques used for machine-generated text detection:
  • Binoculars score (Hans et al., 2024) is a perplexity-based method detailed in Eq. (23) in the Appendix. With this score, generated texts should score lower than human-produced ones. We use the default parameters given by the authors, chosen for their best results on the Binoculars dataset, with falcon-7b-instruct as the detector model, used to calculate perplexity, and falcon-7b as the auxiliary model.
  • DetectGPT (Mitchell et al., 2023) generates minor perturbations of a given text passage, then calculates the differences in log-probabilities between the original text and its perturbed versions. Looking at the curvature of the log-probability function, text from LLMs show less variability and remains in regions of negative curvature when slightly perturbed compared to human-generated text. We used the default parameters provided, with GPT2-XL computing the log probabilities and T5-3b generating the perturbations. To better fit the implementation, we truncated our texts when necessary and selected a random sample of 1,000 pairs of original and generated texts. We run the "100_d" experiment with 100 perturbations applied to each text.
  • Measuring the perplexity (PPL). The text’s log-perplexity as seen by all models in our ensemble –as LLMs usually generate text having a lower perplexity than human’s productions (Vasilatos et al., 2023, Guo et al., 2023, Mitrović et al., 2023, Li et al., 2024)– leads to a very straightforward metric for detecting machine-generated text. PPL (oracle) means we chose the best discriminating (codelength) model, PPL (average) is how good the score is at discriminating with the average of all perplexities.
All the scores reported in this study are ours, using the implementation provided in the original papers.
4 Numerical Results
The following numerical results all use the same configuration, RSA-4, unless explicitly stated otherwise. Our default ensemble of models is composed of TowerBase-7b, TowerBase-13b, Llama-2-7b-chat and Llama-2-7b. All AUC scores are obtained using scikit-learn (Pedregosa et al., 2011).
4.1 Results for the Binocular Benchmark
4.1.1 Baseline ensemble
Our main results are in Table 2, where we report detection scores for the Binoculars datasets (both the original version and our Llama-base addition (results with Mistral-based generations are in Appendix C). For the first, the best two approaches are Binoculars and RSA which outperform the other baselines. Note that contrarily to Binoculars, our results are obtained without any tuning, and without using the generator model in our ensemble. RSA-4 (with averaging, our best variant here) even slightly outperforms this baseline on Pubmed, whose vocabulary is more specialized towards medicine than the one used in CNN and CC News. Results on CNN are harder to interpret, as artificial texts in this dataset appear to be easily uncovered: using a threshold on the perplexity of our models leads to a near-perfect identification. In the following, we ignore this sub-dataset as all methods perform equally well on it.
Pubmed CNN CC_news
Original Regen Original Regen Original Regen
PPL (oracle) 0.823 0.690 0.598 1.000 0.617 0.632
PPL (average) 0.808 0.637 0.566 0.999 0.584 0.569
Detect GPT 0.716 0.668 0.642 0.458 0.580 0.570
Binocular score 0.988 0.633 0.995 1.000 0.979 0.578
RSA-4 (avg) 0.992 0.887 0.993 0.999 0.971 0.854
RSA-4 (max) 0.979 0.904 0.992 0.993 0.948 0.861
RSA-5 (+phi) 0.992 0.885 0.988 0.998 0.950 0.815
Table 2: Detecting artificial texts in the Binoculars Dataset (AUC scores).
4.1.2 Including the generator in our ensemble
For the Llama-7b-based regeneration, we see a large decrease of Binoculars’s performance, as its results become comparable to the PPL thresholding baselines. We note that the (max) variant of robust scoring algorithm becomes the best-performing approach, presumably owing to the strong predictions of the Llama-2 models. If RSA’s results are more stable than Binoculars, the decrease is still significant for Pubmed and CC_news.
This decrease may seem surprising since Llama-2-7b is one of the models we use to detect the generated text. We hypothesise that the sampling method plays here a very important role in the performance of our methods. The Binoculars dataset was generated using a temperature of 0.70.70.70.7777Personal correspondence with the first author., while our regeneration used a temperature of 0.6 along with nucleus sampling ($p=0.9$) and a repetition penalty of 1.18. These parameters are recommended by the official Llama repository [^8] https://github.com/meta-llama/llama/blob/main/example_text_completion.py as they should produce results more diverse, and in a way, human-like than the logits output by raw models. This means that the actual sampling distribution does not fully match the base distribution used in the detector.
In a follow-up experiment, we replace the distribution computed by Llama-7b with a “distorted” version, which approximates the effects of temperature and nucleus-sampling [^9] For this, we first divide the model logits by the temperature, then apply the softmax and perform nucleus filtering with $p=0.9$, making sure to smooth the resulting distribution so that filtered tokens still get a small positive value.. This has a clear effect on the Blahut–Arimoto weights ($q^{*})$, which increase for Llama-7b and decrease for Llama-7b-instruct, but hardly change the detection performance (see Figure 1 in Appendix). This hints at the fact that for generator-detector approaches, the actual generator sampling policy should also be taken into account.
4.1.3 Augmenting the ensemble
RSA makes the augmentation of the ensemble quite easy, as long as all models use the same underlying tokenizer. To showcase the effect of this feature, we add Phi-3 (Abdin et al., 2024) in our ensemble. In this experimental setting, this extension of the model set is of little consequence (see line “+phi” in Table 2), the most notable variation being a small drop in performance for the regenerated CC_news subpart (4% of AUC in the worst-case scenario). Yet, we observe that Phi consistently gets a substantial share of the Blahut–Arimoto weights (see Figure 2(b) in Appendix) and plays a significant role in the classification decision, which has the effect of making Llama-generated texts slightly less likely for the combined model.
4.1.4 Handling poor generators: the case of unigram models
As noted by Hans et al. (2024), simple random generators are hard to detect for Binoculars (as well as for detectors using a PPL-based threshold): this is because random word salads are “surprising” (have a high perplexity) for well-trained detectors, and tend to be confused with human productions. We reproduce this observation using the Random corpus and observe that all baselines, as well as our proposed detector, achieve AUC scores close to 00 for this dataset. The PPL of word salads is however much larger than for human texts, so setting an upper bound of the human PPL would, in that case, provide a very easy fix.
It is tempting to see whether adding such a poor generator into our combination is of any help. For this experiment, we train a unigram language model [^10] We process the Brown Corpus (Francis and Kucera, 1979), tokenized with Llama tools and use count ratios to estimate the probabilities of each token. A smoothing parameter of $\epsilon=1e-10$ is added to each count to avoid null values. and combine it with our four baseline detectors.
The detector obtained with this extended ensemble remains completely unable to sort random from human texts: having the random model in the robust scoring algorithm does not make random texts more likely. This model is also slightly detrimental when it comes to detecting strong generators, as we observe a mild drop in performance throughout (see Appendix, Table 4 for the Binoculars dataset). This is because the unigram model predicts unexpected tokens most of the time and therefore often gets a significant weight in the Blahut–Arimoto weights (see Figure 2(c) in Appendix), which leads to a huge discrepancy between the cross-entropies scores that are averaged in Definition 1.
4.2 Detecting ChatGPT: Ghostbuster and M4
For Ghostbuster, a threshold on the perplexity is enough to achieve an AUC over $0.99$. As a matter of fact, those ChatGPT generations are quite easy to detect. Full results for this dataset are given in Table 7 (Appendix D), where we see that Binoculars gets the highest scores, while RSA-4 and PPL based baselines also give very good AUC scores.
For multilingual data, however, the task is much harder. We considered for this experiment the Arabic, Russian, and Bulgarian datasets from the M4 paper Wang et al. (2024a), as these languages are better represented in the training set of the Tower Family of LLMs than in the Llama-2 family (used in our ensemble) and the Falcon family (used by the Binoculars score). Having multilingual detector models enables us to obtain good performance for Arabic, Bulgarian, and Chinese, less so for Russian. Binoculars scores remain strong for Chinese, but tend to drop for the other languages. When it comes to model selection, our method appears to give more weight to the TowerBase models in our ensemble compared to when just looking at English texts (see Figure 4 in Appendix). This nicely illustrates the benefits of using multiple and complementary models, each one with a restricted domain of “expertise”.
Method Arabic Russian Bulgarian
PPL (oracle) 0.918 0.694 0.932
BinocularsScore 0.686 0.674 0.742
RSA (ours) 0.909 0.737 0.974
Table 3: AUC for Multilingual ChatGPT.
4.3 Detecting scholarly texts
As previously mentioned, our dataset of scholarly texts is composed of fully generated papers. As they were produced by a GPT-2 model fine-tuned on Arxiv’s “computation and language” section, the distribution of tokens greatly differs from all models considered in this study. Therefore, the generated texts are completely out-of-domain and get a high perplexity, which makes them difficult to discriminate from human texts (see Figure 3 in the Appendix) with perplexity-based models. This holds, in particular, for RSA-4 and Binoculars, which achieve respective AUCs of 0.421 and 0.505. This shows that brittleness issues are not fully solved, and hints at including more domain-adapted models in detector ensembles.
5 Related Work
The improvement of text generation abilities of LLMs has raised many concerns about potential misuses such as generating disinformation (Zellers et al., 2019), abusive content (Crothers et al., 2023), forged academic publications (Liu et al., 2024), or to cheat during academic exams (Vasilatos et al., 2023). Such fake texts seem difficult for humans to spot (Gehrmann et al., 2019). In response, the issue of automatically detecting machine-generated texts has been subject to an increasing focus. This problem can be framed in multiple ways: as a binary (human vs. non-human) decision problem, as the problem of detecting one known artificial agent (e.g.,  ChatGPT (Mitrović et al., 2023; Liu et al., 2024)), or as discriminating the correct model in a predefined list (Li et al., 2023). As texts are extremely varied, another distinction is between closed-domain (e.g. scientific (Liyanage et al., 2022), academic (Liu et al., 2024) or user-generated content (Fagni et al., 2021; Kumarage et al., 2023)) vs. open-domain text detection. Assuming the generator models are known, various settings can be considered, depending on whether models can be openly queried (open parameters), whether they expose their full logits, or just the top prediction (and associated probability), etc.
Supervised detection, especially when one single generator is used, looks like a simple problem, as detection rates in the high 90s are easily achieved (Zellers et al., 2019; Guo et al., 2023; Liu et al., 2024) using strong text classifiers based on features extracted by Roberta (Conneau et al., 2020) or T5 (Raffel et al., 2020). These approaches are brittle, though, and their success depends on particular generator-detector pairs (Antoun et al., 2024), prompting e.g. Verma et al. (2023) to design automatic feature extractors from an ensemble of detectors to improve the robustness of their system.
Unsupervised detection is somewhat more challenging. Most approaches to date rest on the intuition that human-written texts are more “surprising” than machine-generated texts [^11] Assuming generation does not use random sampling, in which case the reverse is likely to be observed, as long artificial texts drift away from natural writings (Zellers et al., 2019)., which should translate into a tangible difference in token-wise log-probability121212(Mitchell et al., 2023) argues that the difference is better seen at the level of log-ranks. between the two. This idea is already used in GPTzero [^13] https://gptzero.me/ and thresholding perplexity usually provides strong baselines (see, inter alia, (Gehrmann et al., 2019; Ippolito et al., 2020; Mitchell et al., 2023)). Such techniques heavily rely on the detector model(s) used to compute the log-likelihood of artificial and human-generated texts, which must be robust to variations in text domains, genres, styles, and, needless to say, languages (Wang et al., 2024a); as well as variations in the generator itself Antoun et al. (2024).
Mitchell et al. (2023) and Bao et al. (2024) exploit a similar intuition, arguing that small random perturbations [^14] Performed by a strong generation model, T5 Raffel et al. (2020) in their study. of an artificial text will on average make it less likely (more surprising), which is not true for human-written texts. They derive a statistical criterion to implement this idea, based on an estimate of the curvature of the log-probability function, and report near-perfect detection scores on three types of texts, generated by 5 models. As explained above, Hans et al. (2024) also rely on a function of the per-token log-perplexity, in their case contrasted with the cross-entropy of an auxiliary model.
These valuable works point to the over-reliance on one specific detector model as a major limitation of the state-of-the-art. Our proposed mitigation relies on ensemble techniques, that are also considered in the supervised detection setting, e.g. in (Verma et al., 2023; Wang et al., 2023; El-Sayed and Nasr, 2023; Liyanage and Buscaldi, 2023).
Abandoning generator-detector-based techniques altogether, (Mao et al., 2024; Yang et al., 2024) develop effective detection approaches based on regeneration strategies, prompting the (known) generator to regenerate part of the input text. The intuition is that artificial inputs are likely to be regenerated exactly, while human texts exhibit greater redundancy, resulting in the need for longer codelengths. Other strategies to combat AI-generated text detection have also been considered. Text watermarking (Kirchenbauer et al., 2023a, b; Liu and Bu, 2024), is one of them. Its efficiency and robustness, however, are still subject to discussions, e.g., (Zhang et al., 2023).
Several recent works have focused on the robustness of detection methods. When subjecting machine-generated text to simple modifications, in (Wang et al., 2024b), only watermarking remains able to accurately identify artificial documents. Dugan et al. (2024) present a diverse set of artificial texts, generated with multiple models and sampling strategies, and additionally subject to various adversarial attacks. They also evaluate the robustness of several standard detectors with respect to these source variations, observing that most suffer large drops in performance. In their comparison, Binoculars (Hans et al., 2024) stands out, achieving decent detection scores at False Positive Rates under 1%. Some training strategies have been proposed by Wang et al. (2024b) and Huang et al. (2024) to improve on these robustness issues.
6 Summary and Discussion
As demonstrated, our RSA method is effectively harnessing the strength of the models in the ensemble, offering a scalable solution adaptable to various languages and domains, and also evolving over time. Yet, it displayed varying levels of performance across different text domains, suggesting a need for further research into the domain-adapted models on domain-specific ensemble strategies to improve performance in niche areas (e.g., the academic dataset we used). Furthermore, given that RSA-(max) sometimes outperforms RSA-(avg) and at other times does not, it is crucial to enhance the robustness of our approach. We must develop information-theoretic tools for understanding and selecting the most useful models while filtering out the less effective (or toxic) ones, as demonstrated in the case of the unigram model.
7 Limitations
Our method has proven to be quite effective in detecting content generated by a variety of diverse generators. We however acknowledge a series of shortcomings that will need to be addressed in future work. First, RSA is compute-intensive, as it requires, for each input token, to (a) compute the forward pass for $N$ LLMs, all having significant memory footprints and compute requirements; (b) to run the Blahut-Arimoto algorithm. As a result, each document in our implementation requires 10-20s per document, using a machine equipped with 4 GPUs (V100, 32G). A second limitation of our current approach is the requirement for each detector model to share the same event space, hence the same underlying tokenization. There is no simple solution available in the literature to this problem, which is also encountered in teacher-student knowledge distillation approaches (Raman et al., 2023). Finally, while we have tried to consider multiple generators and datasets for artificial text detection, our benchmarks are still heavily biased toward specific textual genres and languages, in particular clean and relatively long English texts, that correspond to “pure” generations. Testing our system also on other types of data, arguably more challenging, using new benchmarks such as M4GT-bench (Wang et al., 2024c), will help strengthen our robustness claims.
8 Ethics Statement
It should be acknowledged that these tools are not infallible and consequently should not be used as the sole basis for punitive actions or decisions that could affect individuals or organizations. Such methods must be complemented by human oversight and verification before taking any drastic measure to ensure fairness. Moreover, in the course of this work, we have generated medical information and news articles using datasets "Pubmed", "CNN" and "CC_News" to test our method. While we intend to distribute these texts along with our implementation, we must enforce that those are only intended for research use and by no means should be circulated outside of this context, nor be presented as factual content.
Appendix A Proof of Proposition 1
Proof.
We need to show the fundamental identity:
$\Gamma(\mathbf{y}_{<t})$ $\triangleq\min_{q\in\mathcal{P}(\Omega)}\max_{m\in\mathcal{M}}\mathcal{R}_{\theta}(m,q\,;\mathbf{y}_{<t})$ (5)
$=\max_{\mu\in\mathcal{P}(\mathcal{M})}\,\mathcal{I}(\mathbb{M};Y_{t}|\mathbf{y}_{<t}),$ (6)
where the optimal $q^{\star}(y_{t}|\mathbf{y}_{<t})$ achieving the minimum is characterized by the mixture:
$q^{\star}(y_{t}|\mathbf{y}_{<t})=\sum\limits_{m\in\mathcal{M}}\mu^{\star}(m|\mathbf{y}_{<t})p_{\theta}(y_{t}|m,\mathbf{y}_{<t}),$ (7)
and the distribution $\mu^{\star}(m|\mathbf{y}_{<t})$ of the random variable $\mathbb{M}$ on $\mathcal{M}$ follows by solving:
$\mu^{\star}(m|\mathbf{y}_{<t})\triangleq\operatorname*{argmax}_{\mu\in\mathcal{P}(\Omega)}\,\mathcal{I}\big{(}\mathbb{M};Y_{t}|\mathbf{y}_{<t}\big{)}.$ (8)
To this end, we start from the definition $\mathcal{R}_{\theta}$:
$\mathcal{R}_{\theta}(m,q\,;\mathbf{y}_{<t})\triangleq$ $\underset{y_{t}\sim p_{\theta}(y_{t}|m,\mathbf{y}_{<t})}{\mathbb{E}}\left[-\log q(y_{t}|\mathbf{y}_{<t})\right]$
$-\min_{p_{\theta}\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})}\,\underset{y_{t}\sim p_{\theta}(y_{t}|m,\mathbf{y}_{<t})}{\mathbb{E}}\Big{[}-\log p(y_{t}|\mathbf{y}_{<t})\Big{]}$ (9)
$=\underset{y_{t}\sim p_{\theta}(y_{t}|m,\mathbf{y}_{<t})}{\mathbb{E}}\left[-\log q(y_{t}|\mathbf{y}_{<t})\right]-\mathcal{H}_{\theta}(Y_{t}|m,\mathbf{y}_{<t})$ (10)
$=\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(Y_{t}|m,\mathbf{y}_{<t})\big{\|}q(Y_{t}|\mathbf{y}_{<t})\Big{)},$ (11)
where $\mathcal{D}_{\textrm{KL}}(\cdot\|\cdot)$ denotes the Kullback–Leibler divergence. Hence, we can formally state our problem as follows:
$\Gamma(\mathbf{y}_{<t})$ $=\min_{q\in\mathcal{P}(\Omega)}\max_{m\in\mathcal{M}}\mathcal{R}_{\theta}(m,q\,;\mathbf{y}_{<t})$
$=\min_{q\in\mathcal{P}(\Omega)}\max_{m\in\mathcal{M}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$
$=\min_{q\in\mathcal{P}(\Omega)}\max_{\mu\in\mathcal{P}(\mathcal{M})}\underset{m\sim\mu}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)},$ (12)
where the minimum is taken over all the possible distributions $q\in\mathcal{P}(\Omega)$, representing the expected value of regret of $q$ w.r.t. the worst-case distribution over $\mu\in\mathcal{P}(\mathcal{M})$. Notice that this is equivalent to the average worst-case regret Barron et al. (1998); Silva and Piantanida (2022). The equality in (12) holds by noticing that
$\max_{\mu\in\mathcal{P}(\mathcal{M})}\,\underset{m\sim\mu}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$ $\leq\max_{m\in\mathcal{M}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$ (13)
and moreover,
$\max_{m\in\mathcal{M}}\,\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$ $=\underset{m\sim\widetilde{\mu}}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$ (14)
by choosing the measure $\widetilde{\mu}$ to be an uniform probability over the set $\widetilde{\mathcal{M}}$, which is defined as the set of maximizers:
~$\widetilde{\mathcal{M}}=\operatorname*{argmax}_{m\in\mathcal{M}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)},$
and zero otherwise.
The convexity of the KL-divergence allows us to rewrite expression (12) as follows:
$\min_{q\in\mathcal{P}(\Omega)}\max_{\mu\in\mathcal{P}(\mathcal{M})}\underset{m\sim\mu}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}=$ $\max_{\mu\in\mathcal{P}(\mathcal{M})}\min_{q\in\mathcal{P}(\Omega)}\underset{m\sim\mu}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}.$ (15)
This follows by considering a zero-sum game with a concave-convex mapping defined on a product of convex sets. The sets of all probability distributions $\mathcal{P}(\mathcal{M})$ and $\mathcal{P}(\Omega)$ are two nonempty convex sets, bounded and finite-dimensional. On the other hand, $(\mu,q)\rightarrow\underset{m\sim\mu}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$ is a concave-convex mapping, i.e.,
$\mu\rightarrow\underset{m\sim\mu}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$
is concave and,
$q\rightarrow\underset{m\sim\mu}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$
is convex for every $(\mu,q)$, respectively. Then, by classical min-max theorem von Neumann (1928), we have that (15) holds.
Finally, it reminds to show that
$\min_{q\in\mathcal{P}(\Omega)}\underset{m\sim\mu}{\mathbb{E}}$ $\,\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}=\mathcal{I}(\mathbb{M};Y_{t}|\mathbf{y}_{<t}),$ (16)
for any random variable $\mathbb{M}$ distributed according to the probability distribution $\mu\in\mathcal{P}(\mathcal{M})$ and each distribution $p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$.
We begin by showing that
$\underset{m\sim\mu}{\mathbb{E}}$ $\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}\geq\mathcal{I}(\mathbb{M};Y_{t}|\mathbf{y}_{<t})$
for all distributions $q(\cdot|\mathbf{y}_{<t})$ and $p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$. To this end, we consider the following identities:
$\underset{m\sim\mu}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$ $=\underset{m\sim\mu}{\mathbb{E}}\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}p_{\theta}(\cdot|\mathbf{y}_{<t})\Big{)}$
$+\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$
$=\mathcal{I}(\mathbb{M};Y_{t}|\mathbf{y}_{<t})+\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$
$\geq\mathcal{I}(\mathbb{M};Y_{t}|\mathbf{y}_{<t}),$ (17)
where $p_{\theta}(\cdot|\mathbf{y}_{<t})$ denotes the marginal distribution of $p_{\theta}(\cdot|m,\mathbf{y}_{<t})$ w.r.t. $\mu$ and the last inequality follows since the KL divergence is non-negative. Finally, it is easy to check that by selecting:
$q^{\star}(y_{t}|\mathbf{y}_{<t})=\underset{m\sim\mu}{\mathbb{E}}\big{[}p_{\theta}(y_{t}|m,\mathbf{y}_{<t})\big{]}$ (18)
the lower bound in (17) is achieved:
$\min_{q\in\mathcal{P}(\Omega)}\underset{m\sim\mu}{\mathbb{E}}\,\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q(\cdot|\mathbf{y}_{<t})\Big{)}$ $=\underset{m\sim\mu}{\mathbb{E}}\,\mathcal{D}_{\textrm{KL}}\Big{(}p_{\theta}(\cdot|m,\mathbf{y}_{<t})\big{\|}q^{\star}(\cdot|\mathbf{y}_{<t})\Big{)},$ (19)
for every $\mu\in\mathcal{P}(\mathcal{M})$, which proves the identity in expression (16).
The claim in (6) follows by taking the maximum overall probability measures $\mu\in\mathcal{P}(\mathcal{M})$ at both sides of (16), and combining the resulting identity with expressions (15) and (12). The mixture in (8) follows from expression (18) which is a necessary condition to solve the min-max problem. ∎
Appendix B Blahut–Arimoto Algorithm
B.1 Algorithm
Our channel can be specified using two discrete random variables $(\mathbb{M},Y_{t})$ with alphabets $(\mathcal{M},\Omega)$ and probability distributions $\mu$ and $p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$, respectively, conditioned on $\mathbf{y}_{<t}$. The problem to be solved is the maximization of the mutual information:
$\Gamma(\mathbf{y}_{<t})\triangleq\max_{\mu\in\mathcal{P}(\mathcal{M})}\,\mathcal{I}_{\theta}\big{(}\mathbb{M};Y_{t}|\mathbf{y}_{<t}\big{)}.$ (20)
Now if we denote the cardinality $|\mathbb{M}|=M$, $|\Omega|=N$, then $p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$ is an $M\times N$ matrix, which we denote the $i$-th row, $j$-th column entry by $w_{ij}$. For the case of channel capacity, the algorithm was introduced in Arimoto (1972); Blahut (1972) to solve (20). They both found the following expression for the capacity of a discrete channel with channel law $w_{ij}$:
$\Gamma(\mathbf{y}_{<t})=\max_{{\mu}}\max_{Q}\sum_{i=1}^{M}\sum_{j=1}^{N}\mu_{i}w_{ij}\log\left(\frac{q_{ji}}{\mu_{i}}\right),$
where $\mathbf{\mu}$ and $Q$ are maximized over the following requirements:
  • ${\mu}\triangleq(\mu_{1},\dots,\mu_{M})$ is a probability distribution on $\mathcal{M}$. That is, $\sum_{i=1}^{M}\mu_{i}=1$.
  • $Q=(q_{ji})$ is an $N\times M$ matrix that behaves like a transition matrix from $\Omega$ to $\mathcal{M}$ with respect to the channel law. That is, for all $1\leq i\leq M$, $1\leq j\leq N$:
    $q_{ji}\geq 0,\quad q_{ji}=0\Leftrightarrow w_{ij}=0,$
    and every row sums up to 1: $\sum_{i=1}^{M}q_{ji}=1.$
Then, upon initializing a probability measure $\mathbf{\mu}^{0}=(\mu_{1}^{0},\mu_{2}^{0},\ldots,\mu_{M}^{0})$ on $\mathcal{M}$, we can generate a sequence $(\mathbf{\mu}^{0},Q^{0},\mathbf{\mu}^{1},Q^{1},\ldots)$ iteratively as follows:
$(q_{ji}^{t})=\frac{\displaystyle\mu_{i}^{t}w_{ij}}{\displaystyle\sum_{k=1}^{M}\mu_{k}^{t}w_{kj}},$ (21)
and
$\mu_{k}^{t+1}=\frac{\displaystyle\prod_{j=1}^{N}(q_{jk}^{t})^{w_{kj}}}{\displaystyle\sum_{i=1}^{M}\prod_{j=1}^{N}(q_{ji}^{t})^{w_{ij}}},$ (22)
for $t=0,1,2,\ldots$.
Then, using the theory of optimization, specifically coordinate descent, it has been showed that the sequence indeed converges to the required maximum. That is,
$\lim_{t\to\infty}\sum_{i=1}^{M}\sum_{j=1}^{N}\mu_{i}^{t}w_{ij}\log\left(\frac{q_{ji}^{t}}{\mu_{i}^{t}}\right)=\Gamma(\mathbf{y}_{<t}).$
So given a channel law $p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$, the (20) can be numerically estimated up to arbitrary precision.
B.2 Computational complexity
The computational complexity of the Blahut-Arimoto algorithm can be characterized as follows:
  • Number of iterations. The algorithm typically converges linearly, so the number of iterations required, denoted as $T$, is proportional to the desired accuracy of the solution.
  • Operations per iteration. Each iteration involves updating the probability measures in (21) and (22), and evaluating the mutual information, which requires matrix manipulations. Let $M$ and $N$ be the cardinalities of the input and output alphabets, respectively. Each iteration involves operations over all input-output pairs, requiring $\mathcal{O}(M\times N)$ operations.
Combining these, the overall computational complexity of the Blahut-Arimoto algorithm is $\mathcal{O}(T\times n\times m)$, reflecting its dependence on the sizes of $M$ (number of LLMs in the considered family) and $N$ (the vocabulary), and the number of iterations needed for convergence, which depends intrinsically on the underlying distributions.
Appendix C Binocular Results
C.1 Binoculars score
The binoculars score $B(\mathbf{w})$ for an input sequence $\mathbf{w}=\langle w_{0},w_{1},\ldots\rangle$ is defined by
$B(\mathbf{w})\triangleq\frac{\displaystyle\displaystyle\sum_{t=1}^{T}\sum_{y_{t}\in\Omega}\mathbbm{1}[y_{t}=w_{t}]\mathcal{L}_{\theta}(y_{t}|m,\mathbf{y}_{<t})}{\displaystyle\sum_{t=1}^{T}\sum_{y_{t}\in\Omega}\displaystyle p_{\theta}(y_{t}|m^{\prime},\mathbf{w}_{<t})\mathcal{L}_{\theta}(y_{t}|m,\mathbf{y}_{<t})},$ (23)
where the choices of the LLMs $p_{\theta}(\cdot|m,\mathbf{w}_{<t})\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ and $p_{\theta}(\cdot|m^{\prime},\mathbf{w}_{<t})$ $\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ are fundamentally important and have to be empirically selected. Indeed, this represents the main weakness of this score since in practice the best choice of the pair of LLMs $(m,m^{\prime})$ may not be distribution free.
C.2 Numerical results
We further develop on our Binoculars regeneration using Llama-2-7b on table 4, where it can be seen that the log-probabilities of distribution $q^{\star}(4)$ are as good a discriminator of generated texts as the average perplexity obtained with all 4 of our models. We believe this is because Llama-2-7b chat is almost always the best predictor for those samples. This behaviour changes on the Ghostbuster dataset when another instruct model, phi-3-mini-4k-instruct, is added to our ensemble (the reader is referred to 7 for further details). In Table 6, we also report the True Positive Rates at 5% False positive rate. We can see that the perplexities of our models are not good enough discriminators. RSA displays the most robustness across all three generators and domains. Note that in our case, a positive sample is a generated one, as the original Binoculars dataset has more generated samples than human ones, whereas our complementary generations keep the ratio 1:1.
Figure 1 shows the result of the follow-up study mentioned in Section 4.1. When computing the Blahut-Arimoto algorithm on the Pubmed dataset with our default RSA-4 configuration, the weight distribution is as shown on Figure 1(a). We applied a truncated scale to more effectively illustrate the discussed phenomena. This adjustment was necessary due to the high incidence of near-zero values, which would otherwise distort the visual representation of the data. When modifying the Llama-2-7b logits to match the sampling parameters used during generation, the algorithm gives much more importance to the aforementioned model, confirming that this modification suits the distribution of the data better.
Pubmed CNN CC_news
Original Regen-L Original Regen-L Original Regen-L
PPL oracle 0.823 0.690 0.598 1.000 0.617 0.632
PPL average 0.808 0.637 0.566 0.999 0.584 0.569
$q^{\star}$ (4) log-probs 0.807 0.641 0.566 1.000 0.579 0.568
Binocular score 0.988 0.633 0.995 1.000 0.979 0.578
RSA-4 (avg) 0.992 0.887 0.993 0.999 0.971 0.854
RSA-4 (max) 0.979 0.904 0.992 0.993 0.948 0.861
RSA-5 (+phi) 0.992 0.885 0.988 0.998 0.950 0.815
RSA-5 (+unigram) 0.986 0.846 0.960 1.000 0.898 0.766
Table 4: More results the Binoculars Dataset (AUC scores).
Pubmed CNN CC_news Average
Original Regen-L Regen-M Original Regen-L Regen-M Original Regen-L Regen-M
PPL TowerBase-7b 0.798 0.631 0.806 0.557 0.999 0.639 0.582 0.573 0.669 0.695
PPL TowerBase-13b 0.823 0.528 0.801 0.561 0.999 0.641 0.578 0.439 0.652 0.669
PPL Llama-2-7b-chat 0.817 0.690 0.843 0.598 1.000 0.716 0.617 0.632 0.718 0.737
PPL Llama-2-7b 0.786 0.676 0.820 0.544 1.000 0.704 0.556 0.617 0.693 0.711
$q^{\star}$ log-probs 0.807 0.641 0.836 0.566 1.000 0.727 0.579 0.568 0.703 0.714
Binocular score 0.988 0.633 0.892 0.995 1.000 0.956 0.979 0.578 0.922 0.883
RSA-4 (avg) 0.992 0.887 0.961 0.993 0.999 0.971 0.971 0.854 0.940 0.952
RSA-4 (max) 0.979 0.904 0.960 0.992 0.993 0.987 0.948 0.861 0.943 0.952
Table 5: Comparing results on Mistral-7B (Regen-M) and Llama-2-7B (Regen-L) generations (AUC scores).
Pubmed CNN CC_news Average
Original Regen-L Regen-M Original Regen-L Regen-M Original Regen-L Regen-M
PPL TowerBase-7b 0.318 0.216 0.476 0.051 0.998 0.106 0.204 0.026 0.178 0.286
PPL TowerBase-13b 0.365 0.099 0.480 0.066 0.999 0.109 0.182 0.016 0.174 0.277
PPL Llama-2-7b-chat 0.353 0.297 0.546 0.060 1.000 0.183 0.230 0.027 0.215 0.323
PPL Llama-2-7b 0.293 0.279 0.500 0.049 1.000 0.173 0.187 0.033 0.227 0.305
$q^{\star}$ log-probs 0.330 0.206 0.533 0.059 0.999 0.191 0.195 0.024 0.245 0.309
Binocular score 0.942 0.278 0.683 0.975 1.000 0.835 0.924 0.195 0.561 0.710
RSA-4 (avg) 0.963 0.615 0.814 0.971 1.000 0.898 0.868 0.552 0.730 0.823
RSA-4 (max) 0.904 0.660 0.800 0.964 0.967 0.954 0.782 0.538 0.742 0.812
Table 6: Comparing results on Mistral-7B (Regen-M) and Llama-2-7B (Regen-L) generations (TPR@5%FPR).
(a) Blahut–Arimoto weights without sampling on Llama-2-7b logits.
(b) Blahut–Arimoto weights with sampling on Llama-2-7b logits.
Figure 1: Comparison of Blahut–Arimoto weights with and without sampling on Llama-2-7b logits when looking at text generated with the same parameters.
(a) Blahut–Arimoto weights on Pubmed generated with Llama-2-7b.
(b) Blahut–Arimoto weights on Pubmed generated with Llama-2-7b with Phi in the ensemble.
(c) Blahut–Arimoto weights on Pubmed generated with Llama-2-7b with our Unigram model in the ensemble
Figure 2: Comparison of Blahut–Arimoto weights on the Pubmed Regen-L dataset for different RSA  configurations.
Appendix D Ghostbuster Results
When it comes to the Ghostbuster dataset, all methods perform extremely well, with RSA-4 lagging a bit behind. However, unlike for Binoculars, adding Phi into the mix is of great help. We believe this is because ChatGPT’s distribution is closer to Phi’s than the other models in our mix. Adding one more model to our RSA-4 thus enables us to approach the generator’s distribution better.
Method Writing Prompts Reuter Essay Average
PPL TowerBase-7b 0.973 0.961 0.990 0.975
PPL TowerBase-13b 0.976 0.964 0.991 0.977
PPL Llama-2-7b-chat 0.989 0.994 0.994 0.992
PPL Llama-2-7b 0.977 0.971 0.991 0.980
PPL Phi-3-mini-4k-instruct 0.993 0.990 0.995 0.993
$q^{\star}$ log-probs 0.985 0.984 0.994 0.988
$q^{\star}$ 5 (+phi) log-probs 0.992 0.990 0.995 0.992
Binoculars Score 0.990 0.993 0.996 0.993
RSA(avg) 0.946 0.971 0.911 0.943
RSA(max) 0.975 0.979 0.974 0.976
RSA-5 (+phi, avg) 0.975 0.986 0.966 0.976
RSA-5 (+phi, max) 0.993 0.996 0.995 0.995
Table 7: AUC Scores on the Ghostbuster dataset.
Method Writing Prompts Reuter Essay Average
PPL TowerBase-7b 0.910 0.789 0.961 0.887
PPL TowerBase-13b 0.924 0.810 0.971 0.902
PPL Llama-2-7b-chat 0.967 0.983 0.977 0.976
PPL Llama-2-7b 0.924 0.854 0.968 0.915
PPL Phi-3-mini-4k-instruct 0.974 0.962 0.987 0.974
$q^{\star}$ log-probs 0.953 0.922 0.982 0.952
$q^{\star}$ 5 (+phi) log-probs 0.974 0.955 0.987 0.972
Binoculars Score 0.960 0.963 0.987 0.970
RSA(avg) 0.737 0.843 0.558 0.713
RSA(max) 0.887 0.886 0.882 0.885
RSA-5 (+phi, avg) 0.899 0.933 0.832 0.888
RSA-5 (+phi, max) 0.979 0.983 0.986 0.983
Table 8: TPR@5%FPR on the Ghostbuster dataset.
Appendix E arxiv-cs.cl
As mentioned in Section 4.3, both the baseline Binoculars and our RSA method perform poorly on this dataset and are unable to effectively differentiate between generated and original texts. As illustrated in Figure 3, the scores computed by both methods on human and artificial texts largely overlap, resulting in a detection performance that is close to a random decision.
Figure 3: Histograms of scores on the scholarly dataset.
Appendix F M4
Table 9 displays all our results on the M4 multilingual datasets we used. We notice that RSAperforms better on Chinese Davinci than Chinese ChatGPT but it is the opposite when it comes to Russian and Bulgarian. On those two languages, performances quickly drop when looking at the TPR@5%FPR.
On Figure 4, we compare the Arimoto weights on texts from the same domain (news articles) but different languages. The English one favours Llama-2-7b very heavily, while the bulgarian version gives more credit to the TowerBase models we use, which is what we expect since those models have seen more multilingual data during training.
AUC Arabic German Chinese Russian Bulgarian
Chatgpt ChatGPT ChatGPT DaVinci ChatGPT DaVinci ChatGPT DaVinci
PPL (oracle) 0.918 0.693 0.983 0.946 0.695 0.391 0.932 0.449
BinocularsScore 0.686 0.914 0.982 0.979 0.674 0.554 0.742 0.659
RSA(avg) 0.909 0.890 0.930 0.980 0.737 0.324 0.974 0.556
RSA(max) 0.912 0.885 0.909 0.964 0.688 0.259 0.956 0.642
TPR @ 5% FPR
PPL (oracle) 0.587 0.052 0.945 0.862 0.199 0.011 0.719 0.016
BinocularsScore 0.129 0.679 0.929 0.957 0.085 0.038 0.202 0.117
RSA(avg) 0.564 0.603 0.609 0.951 0.220 0.016 0.898 0.064
RSA(max) 0.558 0.571 0.498 0.864 0.139 0.013 0.789 0.099
Table 9: Detailed results for the M4 dataset.
(a) Blahut–Arimoto weights when looking at the llama-2-7b generated cc-news dataset.
(b) Blahut–Arimoto weights when looking at Bulgarian ChatGPT generations.
Figure 4: Comparison of Blahut–Arimoto weights between English and Bulgarian.